Questions 5 and 6

Jordan National Bank is considering adding 5 new ATM machines. Each machine costs $35,000 and installation costs are $15,000 per machine. Officials at the bank expect the new machines to save $.33 per transaction and expects 250,000 transactions per year on the new machines. They expect the new machines to last for 15 years.

5. What is the residual value of the machines?

            A. It will be based on the Einstein Residual Factor.

            B. The problem does not stipulate a residual value.

            C. residual value is based on the revenue for year 16.

6. If a return of 9% is required, what is the net present value of this investment?

            A. $929,252

            B. $665,007

            C. $415,000

            D. $306,729

            E. $    2,691

Please show me the work so I can do it on the test. Thank you
Answer #1

Answer 5 .

The problem doesnot stipulate residual value

Answer 6. C =415000

Cost of machine = 5*(35000+15000) =250000

Saving per year = 250000*.33= 82500

NPV = (250000) + 82500/(1+.09)^1+82500(1+.09)^2+...+82500(1+.09)^15 = $415000